OpenStudy (anonymous):

what will be the pH of solution when two buffer of pH 4 and 6 mixed each other.

OpenStudy (trancenova):

Well, if pH 7 is neutral, and 6 and 4 are both equally apart from 7 (and assuming you add the same amount of each buffer), what do you think? :)

OpenStudy (anonymous):

1) \[10^{-4}+10^{-6} = 1.01*10^{-4}\]

OpenStudy (anonymous):

your volume doubled, therefore, (1.01*10^{-4})/2 = 5.05*10^{-5}

OpenStudy (anonymous):

pH = -log (5.05 x 10^-5) = 4.30

OpenStudy (anonymous):

I think you can not tell just by the facts you gave us. for example: if the buffer capacity of the pH 6 solution is used up, but the pH 4 buffer is still able to compensate for the change, the pH will be close to 4, and vice versa. you need to give concentrations of HA / A- for each buffer solution.

OpenStudy (anonymous):

lol @ trancenova, btw =) 6 and 4 being equally apart from 7, nice

OpenStudy (anonymous):

@Schleifspur: yes, thank you, my mistake, I assumed that the volumes of the buffers would be the same, but that is not clear from the question. Thanks for catching!!!!

OpenStudy (blues):

You'd need to know the concentrations of both buffers.

OpenStudy (anonymous):

@ blues: pH already shows molar concentration of H30+ ions

OpenStudy (blues):

Oh, didn't notice that.